OpenStudy (aihberkhan):

Question below:

OpenStudy (aihberkhan):

Is \(\frac{ 3 }{ 10 } * 10\) the same thing as \(\frac{ 3 }{ 10 } * \frac{ 10 }{ 10 }\)?

OpenStudy (aihberkhan):

@imqwerty

OpenStudy (aihberkhan):

@Michele_Laino

Mehek (mehek14):

no \(\dfrac{3}{10}*10=\dfrac{3}{10}*\dfrac{10}{1}\)

OpenStudy (aihberkhan):

Oh I see :) Thanks. I knew it was different but I just couldn't get my head straight :) Thanks! :) @Mehek14

OpenStudy (skullpatrol):

10 divided by 10 equals 1

OpenStudy (skullpatrol):

Any nonzero divided by itself is 1. That is what is meant by $$\huge a\times\dfrac{1}{a}=1$$ For a not equal to 0